# ScholaWrite: A Dataset of End-to-End Scholarly Writing Process

### v3

@@ 5.1 Predicting Next Writing Intention
Iter. | Llama-8B-SW | Llama-8B-Zero | GPT-4o |
[…]
Results
Table 3 presents the weighted F1 scores for predicting writing intentions across baselines and fine-tuned models. Regardless of the intricate nature of the task itself666Each model predicts the next intention using only the ”before” text, while human annotators consider multiple ”before and after” edits. Moreover, the chosen next intention is not necessarily the only correct one., all models finetuned on ScholaWrite show an improved performance compared to their baselines. BERT and RoBERTa achieved the most improvement, while LLama-8B-Instruct showed a modest improvement after fine-tuning. As detailed in Appendix E, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs. This aligns with findings from Grasso and Locci (2024); Yu et al. (2023), which shows that RoBERTa and BERT can often match or even outperform LLMs for the text classification tasks. Those results demonstrate the effectiveness of our ScholaWrite dataset to align language models with writers’ intentions.
5.2 Iterative Self-Writing
Setups
During iterative self-writing (Figure 4), a fine-tuned model processes LaTeX-formatted seed document (as “before-text”) with a context prompt to predict the next intention, then revises the text (“after-text”) accordingly given prompt. The revised document then serves as the new seed for the next iteration. This process repeats until a set iteration limit (e.g., 100) is reached. All models use the same train (80%)-test (20%) split across experiments. See Appendix E and Figure 14 for training details.
We fine-tune Llama3.1-8B-Instruct (Llama-8B-SW) and compare it to vanilla Llama-8B-Instruct (Llama-8B-Zero) and GPT-4o. Also, seed documents were derived from LaTeX-formatted abstracts of four award-winning NLP papers on diverse topics Zeng et al. (2024); Lu et al. (2024b); Du et al. (2022a); Etxaniz et al. (2024) (Appendix LABEL:table:seed-entry-johnny-LABEL:table:seed-entry-latxa).
Metrics
We evaluated lexical diversity (unique-to-total token ratio), topic consistency (cosine similarity between seed and final output), and intention coverage (unique writing intentions used over 100 iterations). For human evaluation, three native English speakers with LaTeX expertise assessed outputs from Llama-8B-Zero and Llama-8B-SW on accuracy (alignment with predicted intention), alignment (similarity to human writing), fluency(grammatical correctness), coherence(logical structure), and relevance (connection to the seed document) - Refer to Appendix LABEL:sec:appendix:human-eval. Accuracy was judged per iteration, while other metrics used pairwise comparisons. Inter-annotator agreement (IAA)777The IAA scores are 0.84 (SW) and 0.76 (Zero) for accuracy, all 100% for alignment, fluency, coherence, and 49.8% (SW) and 100% (Zero) for relevance. was measured using Krippendorff’s alpha for accuracy and percentage agreement for others.
Metrics | Model | Seed 1 | Seed 2 | Seed 3 | Seed 4 |
Accuracy | SW | 21.0 | 10.3 | 18.3 | 15.3 |
Zero | 35.7 | 29.7 | 45.3 | 43.3 | |
Alignment | SW | 0 | 0 | 0 | 0 |
Zero | 3 | 3 | 3 | 3 | |
Fluency | SW | 0 | 0 | 0 | 0 |
Zero | 3 | 3 | 3 | 3 | |
Coherence | SW | 0 | 0 | 0 | 0 |
Zero | 3 | 3 | 3 | 3 | |
Relevance | SW | 1 | 3 | 2 | 1 |
Zero | 3 | 3 | 3 | 3 |
Results
However, our human evaluation (Table 5) revealed that Llama-8B-SW generated less human-like writing, in terms of fluency and logical claims. It also struggled with generating texts aligned with the predicted intentions. See Appendix Tables LABEL:table:human-eval-seed1 to LABEL:table:human-eval-seed4 for more details. Despite the weaknesses, Llama-8B-SW still produced more relevant content (Seed 2), which aligns with topic consistency trends in Figure 5, highlighting the usefulness of ScholaWrite dataset in certain contexts.
Moreover, Llama-8B-SW exhibited the most human-like writing activity patterns over time (Figure LABEL:fig:writing-step-intention-all-model), which frequently switches between implementation and revision and covers all three high-level processes. Llama-8B-Zero and GPT-4o tend to remain in a single high-level stage throughout all 100 iterations of self-writing. Compared to Appendix Figure 12, which depicts frequent transitions across all three stages in an early draft (e.g., the first 100 steps), Llama-8B-SW most closely replicates human writing behaviors in iterative writing tasks. These findings reinforce the potential of ScholaWrite in helping LLMs emulate human scholarly writing processes.
@@ 6 Conclusion and Future Work
We present ScholaWrite, a first-of-its-kind dataset capturing the end-to-end cognitive process of scholarly writing, comprising nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension. This dataset, sourced from ten graduate students with varying levels of scientific writing expertise, is further enriched with expert annotations based on a novel taxonomy of cognitive writing intentions inspired by Flower and Hayes (1981); Koo et al. (2023). Through several experiments, ScholaWrite shows its value for advancing the cognitive capabilities of LLMs and developing cognitively-aligned writing assistants, enabling them to mimic the complex, non-linear, and intention-driven nature of human writing.
Future work includes expanding the dataset to diverse academic fields, authors, and collaborative projects, thus enabling models to generalize beyond fact recall to emulate human decision-making and reasoning in more realistic academic environments. Additionally, integrating advanced memory architectures and lifelong learning techniques could further enhance LLMs’ ability to adapt dynamically to evolving writing intentions and produce coherent, high-quality scholarly outputs.
@@
Furthermore, all participants were early-career researchers (e.g., PhD students) at an R1 university in the United States. Expanding the dataset to include senior researchers, such as post-doctoral fellows and professors, could offer valuable insights into how writing strategies and revision behaviors evolve with research experience and expertise. Despite these limitations, our study captured an end-to-end writing process for 10 unique authors, resulting in a diverse range of writing styles and revision patterns. The dataset contains approximately 62,000 keystrokes, offering fine-grained insights into the human writing process, including detailed editing and drafting actions over time. While the number of articles is limited, the granularity and volume of the data provide a rich resource for understanding writing behaviors. Prior research has shown that detailed keystroke logs, even from small datasets, can effectively model writing processes Leijten and Van Waes (2013); Guo et al. (2018); Vandermeulen et al. (2023). Unlike studies focused on final outputs, our dataset enables a process-oriented analysis, emphasizing the cognitive and behavioral patterns underlying scholarly writing.
@@
Fourth, due to computational and cost constraints, we evaluated the usability of the ScholaWrite dataset with a limited number of LLMs and hyperparameter configurations. As shown in Table 3, the Llama-8B-Instruct model demonstrated only marginal improvements after fine-tuning on our dataset. This underscores the need for future research to explore advanced techniques, such as fine-grained prompt engineering, to better align LLM outputs with human writing processes. Specifically, optimizing prompts with clearer contextual guidance (e.g., "before-text" and intention label definitions) may significantly enhance model performance.
@@ A.2 Technical Details of System Implementations
When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel888To prevent privacy concerns, the extension filters out keystroke data from any unauthorized Overleaf projects. Please see Appendix A.2 for more details.. When each of these actions999Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page. occurs, the extension uses ‘diff_match_patch’ package101010https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 6). Then, the extension will send the array along with metadata (e.g., time stamp, author ID, etc.) to the backend server.
@@ Appendix D ScholaWrite dataset statistics
Table 6 presents the overall statistics of our ScholaWrite dataset. Table 8 shows the distribution of intention labels per Overleaf project from ScholaWrite.
Project | 1 | 2 | 3 | 4 | 5 |
# Authors | 1 (3) | 1 (4) | 1 (3) | 1 (4) | 9 (18) |
# keystrokes | 14,217 | 5,059 | 6,641 | 8,348 | 27,239 |
# Words added | 17,387 | 23,835 | 7,779 | 12,448 | 57,511 |
# Words deleted | 11,739 | 15,158 | 2,308 | 7,621 | 25,853 |
@@
Figures 9 to 12 show several characteristics of human writing process, analyzed from ScholaWrite dataset: (1) Figure 9 - the average Wasserstein distance between each intention distribution and uniform distribution; (2) Figure 10 - distribution of labels over time; (3) Figure 11 for high-level intention distribution over time; and (4) Figure 12 for intention-wise writing activity distribution over time.
Appendix E More about ScholaWrite Evaluation
E.1 Training environments
@@
For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct111111https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library121212https://github.com/unslothai/unsloth.
@@
For the ‘after’ text generation subtask from the iterative self-writing experiment (Sec. 5.2), we used the following hyperparameter setups for the fine-tuned Llama-8B-Instruct: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 10; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. Also, it took approximately 12 hours on one NVIDIA L40s.
@@
We used the GPT-4o-2024-08-06 version for the classification and iterative writing experiments. The iterative writing with 100 iterations took approximately 1 hour on each seed.
E.2 Details About Finetuning Process
Next Intention Prediction (Sec 5.1)
The fine-tuning prompt included all possible labels with definitions, task instructions, the “before-text” chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”. Differences in prompts were limited to only task instructions (see §E.3 for prompt details).
To achieve optimal performance while minimizing memory usage, we employed QLoRA Dettmers et al. (2024) to fine-tune all linear modules of a 4-bit quantized Llama. During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s. This ensures the model is trained exclusively on the response portion of the fine-tuning prompt (i.e., the predicted intention label), without being influenced by the instructional components of the input. The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
Iterative Self-Writing (Sec 5.2)
The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. For fine-tuning, we used QLoRA Dettmers et al. (2024) to optimize all linear modules of a 4-bit quantized model while maintaining a small memory footprint. Additionally, the embed_tokens and lm_head modules were set as trainable and saved in the final checkpoint. To focus training on the response portion (the “after” text), we used the train_on_response_only function, masking the task instructions, verbalizer, and “before” text with -100s. This ensures the model learns to generate the "after" text without being influenced by instructional input. The model was trained for one epoch with a batch size of 1, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
E.3 Prompt Templates
E.3.1 Prediction Prompt for Llama-8B-Zero models
@@
E.3.2 Generation Prompt for Llama-8B-Zero models
## Input: before_text ”
E.3.3 Prediction Prompt for Llama-8B-SW models
@@
E.3.4 Generation Prompt for Llama-8B-Zero models
## Input: before_text ”
E.3.5 Classification Prompt for GPT-4o
@@
E.3.6 Generation Prompt for GPT-4o

### v4

@@ 5.3 Are writing patterns stable throughout the process, or do they change across phases?
6 Can LLMs Support Human Writing?
Scientific writing is deeply cognitive, non-linear, and phase-dependent. Writers continuously shift between intents, and multitasking across intertwined intentions. For LLMs to genuinely assist this process, they must (i) infer what the writer is doing or about to do, and (ii) provide support that aligns with the writer’s actual cognitive intent. We evaluate these capabilities through two tests using ScholaWrite:
6.1 Predicting Next Writing Intention
Because writing involves frequent task-switching, an effective assistant should anticipate the writer’s next move. We test whether models can predict the upcoming writing intention based on the current text state.
Setup.
We use the “before-text” from each keystroke pair and prompt the model to predict the next writing intention. Models evaluated include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), Llama3.1-8B-Instruct (Dubey et al., 2024), and GPT-4o, with fine-tuned versions train on our dataset. Performance is measured using weighted F1 to account for label imbalance. See details in Appendix G.2.
Findings.
Base models perform poorly (F1 0.12), indicating that current LLMs cannot infer writing intent from text alone. Fine-tuning on ScholaWrite improves performance dramatically (to 0.64 for BERT/RoBERTa), showing that cognitively annotated writing data helps models anticipate human writing behavior, a key step toward adaptive, context-aware assistance. Training details and explanations for differences across models are reported in Appendix F.1.
6.2 Output Alignment across Writing Sessions
To assist writers effectively, a model must not only infer their intentions but also generate text aligned with those goals. We evaluate LLMs under two settings: (1) single-intention sessions, focused on one cognitive goal, and (2) multi-intention sessions, where multiple intentions intertwine.
Setup.
For each session, models receive the “before-text” and either a single intention label or the sequence of intentions observed in that session, then generate the corresponding “after-text.” We use GPT-5 (OpenAI, 2025) as the base model and evaluate alignment with human writing using BERTScore-F1 (Zhang et al., 2019). Results for other models (e.g., Qwen (Yang et al., 2024)) and metrics (e.g., Levenshtein distance) appear in the Appendix F.2 and show consistent patterns.
Findings.
Model alignment decreases as sessions grow longer or cognitively more complex (Figure 5). As shown in Figure 6, GPT-5 (and Qwen) perform best on Revision intentions (e.g., Clarity, Fluency) but struggle with Planning and Implementation, which demand deeper reasoning and structural organization. Alignment further drops when multiple intentions co-occur within a session, indicating models’ limited ability to integrate overlapping cognitive goals. Overall, current LLMs excel at surface refinement but remain weak at sustaining the intertwined reasoning processes that drive human scholarly writing.
@@ 7 Conclusion and Future Work
This work introduces ScholaWrite, the first dataset to capture the end-to-end cognitive process of scholarly writing. Unlike prior work that centers only on final manuscripts, ScholaWrite reveals how writing unfolds through iterative, multi-intent, and phase-dependent behaviors. Our analyses uncover three key insights. First, scholarly writing is highly non-linear: authors continually alternate between planning, implementing, and revising, rather than progressing in discrete stages. Second, writing intentions evolve across phases, with early stages emphasizing idea development and late stages dominated by refinement and multitasking. Third, current LLMs, though proficient at surface-level revision, struggle to anticipate or integrate complex, cognitively demanding intentions, revealing a critical gap between human writing cognition and model capabilities.
ScholaWrite contributes: (1) a fine-grained lens on the cognitive dynamics of scholarly writing; (2) a benchmark for evaluating LLMs’ alignment with human writing intentions; and (3) a training resource for developing models that adapt to evolving, multi-intent writing behavior. Our Chrome extension enables unobtrusive, in-situ data capture in Overleaf, demonstrating a scalable and ethical method for collecting realistic writing data.
Future work will extend beyond keystroke logs to include pre-writing ideation, collaboration, and multimodal composition (e.g., figure and data integration). By tracing the full scholarly writing lifecycle, we aim to deepen our understanding of human cognitive patterns and to design AI writing assistants that not only complete text, but truly complement how humans think and write.
@@
Furthermore, all participants were early-career researchers (e.g., PhD students) at an R1 university in the United States. Expanding the dataset to include senior researchers, such as post-doctoral fellows and professors, could offer valuable insights into how writing strategies and revision behaviors evolve with research experience and expertise. Despite these limitations, our study captured an end-to-end writing process for 10 unique authors, resulting in a diverse range of writing styles and revision patterns. The dataset contains approximately 62,000 keystrokes, offering fine-grained insights into the human writing process, including detailed editing and drafting actions over time. While the number of articles is limited, the granularity and volume of the data provide a rich resource for understanding writing behaviors. Prior research has shown that detailed keystroke logs, even from small datasets, can effectively model writing processes (Leijten and Van Waes, 2013; Guo et al., 2018; Vandermeulen et al., 2023). Unlike studies focused on final outputs, our dataset enables a process-oriented analysis, emphasizing the cognitive and behavioral patterns underlying scholarly writing.
@@ A.2 Technical Details of System Implementations
When a key-up event fires in a browser, the extension collects the writer’s viewable texts in the code editor panel111To prevent privacy concerns, the extension filters out keystroke data from any unauthorized Overleaf projects. Please see Appendix A.2 for more details.. When each of these actions222Example actions are (1) inserting a space/newline; (2) copy/paste; (3) undo/redo; (4) switching files and (5) scrolling a page. occurs, the extension uses ‘diff_match_patch’ package333https://github.com/google/diff-match-patch to generate an array of differences between two subsequent texts (i.e., Figure 7). Then, the extension will send the array along with metadata (e.g., time stamp, author ID, etc.) to the backend server.
@@ Appendix D ScholaWrite dataset statistics
Table 8 shows the distribution of intention labels per Overleaf project from ScholaWrite.
@@
Figures 10 to 13 show several characteristics of human writing process, analyzed from ScholaWrite dataset: (1) Figure 10 - the average Wasserstein distance between each intention distribution and uniform distribution; (2) Figure 11 - distribution of labels over time; (3) Figure 12 for high-level intention distribution over time; and (4) Figure 13 for intention-wise writing activity distribution over time.
Appendix E More about Analyses
Figure 15 shows transition probability between pair of intentions in early stage and later stage.
Table 9 shows top 6-gram writing intention sequences by session coverage (%)
| 6-gram Sequence | Percentage |
| Text Production Clarity Text Production Clarity Text Production Clarity | 3.22 |
| Clarity Text Production Clarity Text Production Clarity Text Production | 3.22 |
| Text Production Object Insertion Text Production Object Insertion Text Production Object Insertion | 2.34 |
| Object Insertion Text Production Object Insertion Text Production Object Insertion Text Production | 2.34 |
| Idea Generation Idea Organization Idea Generation Idea Organization Idea Generation Idea Organization | 2.05 |
Appendix F More about Experiments
F.1 Predicting Next Writing Intention
F.1.1 Training environments
@@
For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct444https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library555https://github.com/unslothai/unsloth.
@@
We used the GPT-4o-2024-08-06 version.
F.1.2 Details About Finetuning Process
The fine-tuning prompt included all possible labels with definitions, task instructions, the “before-text” chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”. Differences in prompts were limited to only task instructions (see §G.3 for prompt details).
To achieve optimal performance while minimizing memory usage, we employed QLoRA (Dettmers et al., 2024) to fine-tune all linear modules of a 4-bit quantized Llama. During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s. This ensures the model is trained exclusively on the response portion of the fine-tuning prompt (i.e., the predicted intention label), without being influenced by the instructional components of the input. The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
F.1.3 Prompt Templates
Prediction Prompt for Llama-8B-Zero models
@@ F.1.3 Prompt Templates
Prediction Prompt for Llama-8B-SW models
@@
Classification Prompt for GPT-4o
@@
F.1.4 Discussion about Next Intention Prediction Results
According to Table 6, none of them is reaching 0.7 in F1. This is likely due to the intricate nature of the task. The model is asked to predict the next intention by only giving the before text. In the annotation task, the annotator labels the data by looking through multiple consecutive before and after text pairs to determine the current intention rather than looking at the current text to predict the next intention. Another reason is that the next intention chosen by the author does not necessarily mean that it is the only correct intention. We also noticed that BERT and RoBERTa perform much better than the vanilla Llama-8b-Instruct. This is likely due to Llama 8b being under-trained as they have a larger size of parameters, while we fine-tuned it on our data for only one epoch.
F.2 Output Alignment across Writing Sessions
F.2.1 Prompt Templates
Single-intention Session Output Generation
You are an expert writing assistant.
Your role is to improve user-provided text according to specific criteria.
The input will be LaTeX text.
Always output only the improved LaTeX text, with no explanations or notes.
Improve the following LaTeX text according to this criteria:
- {intention}: {intention definition}
Text to improve:
{before_text}
Multi-intention Session Output Generation
You are an expert writing assistant.
Your role is to improve user-provided text according to specific criteria.
The input will be LaTeX text.
Always output only the improved LaTeX text, with no explanations or notes.
Improve the following LaTeX text according to these criteria:
{list of intentions and their definitions}
Text to improve:
{before_text}
F.2.2 Results
Figure 16 shows the relationship between a single-intention session duration and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B.
Figure 17 shows the relationship between a single-intention session duration and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B across different intention categories.
Figure 18 shows the relationship between number of writing intentions in a multi-intention session and alignment score (either Levenshtein ratio or BERTScore-F1) for two models, GPT-5 and Qwen-2.5-14B.
Appendix G More about ScholaWrite Evaluation
We envision ScholaWrite as a valuable resource for training language models and improving future writing assistants for scholarly writing. To evaluate its usability, we conducted experiments training LLMs to mimic the complex, non-linear writing processes of human scholars.
Iteratively generating scholarly writing actions from scratch (called Iterative Self-Writing), mirroring the human writing process: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
Setups
During iterative self-writing (Figure 19), a fine-tuned model processes LaTeX-formatted seed document (as “before-text”) with a context prompt to predict the next intention, then revises the text (“after-text”) accordingly given prompt. The revised document then serves as the new seed for the next iteration. This process repeats until a set iteration limit (e.g., 100) is reached. All models use the same train (80%)-test (20%) split across experiments. Figure 23 for training details.
We fine-tune Llama3.1-8B-Instruct (Llama-8B-SW) and compare it to vanilla Llama-8B-Instruct (Llama-8B-Zero) and GPT-4o. Also, seed documents were derived from LaTeX-formatted abstracts of four award-winning NLP papers on diverse topics (Zeng et al., 2024; Lu et al., 2024; Du et al., 2022a; Etxaniz et al., 2024), as shown in Appendix Listings LABEL:table:seed-entry-johnny-LABEL:table:seed-entry-latxa.
Metrics
We evaluated lexical diversity (unique-to-total token ratio), topic consistency (cosine similarity between seed and final output), and intention coverage (unique writing intentions used over 100 iterations). For human evaluation, three native English speakers with LaTeX expertise assessed outputs from Llama-8B-Zero and Llama-8B-SW on accuracy (alignment with predicted intention), alignment (similarity to human writing), fluency(grammatical correctness), coherence(logical structure), and relevance (connection to the seed document) - Refer to Appendix LABEL:sec:appendix:human-eval. Accuracy was judged per iteration, while other metrics used pairwise comparisons. Inter-annotator agreement (IAA)666The IAA scores are 0.84 (SW) and 0.76 (Zero) for accuracy, all 100% for alignment, fluency, coherence, and 49.8% (SW) and 100% (Zero) for relevance. was measured using Krippendorff’s alpha for accuracy and percentage agreement for others.
Llama-8B-SW
Llama-8B-Zero
GPT-4o
| Iter. | Llama-8B-SW | Llama-8B-Zero | GPT-4o |
[…]
Results
However, our human evaluation (Figure 21) revealed that Llama-8B-SW generated less human-like writing, in terms of fluency and logical claims. It also struggled with generating texts aligned with the predicted intentions. See Appendix Tables LABEL:table:human-eval-seed1 to LABEL:table:human-eval-seed4 for more details. Despite the weaknesses, Llama-8B-SW still produced more relevant content (Seed 2), which aligns with topic consistency trends in Figure 20, highlighting the usefulness of ScholaWrite dataset in certain contexts.
Moreover, Llama-8B-SW exhibited the most human-like writing activity patterns over time (Figure 22), which frequently switches between implementation and revision and covers all three high-level processes. Llama-8B-Zero and GPT-4o tend to remain in a single high-level stage throughout all 100 iterations of self-writing (see Appendix LABEL:fig:writing-step-broad-all-model and LABEL:fig:writing-step-intention-all-model for details). Compared to Appendix Figure 13, which depicts frequent transitions across all three stages in an early draft (e.g., the first 100 steps), Llama-8B-SW most closely replicates human writing behaviors in iterative writing tasks. These findings reinforce the potential of ScholaWrite in helping LLMs emulate human scholarly writing processes.
G.1 Training environments
BERT & RoBERTa
Llama
For all experiments, we used baseline models of 4-bit quantized Llama-8B-Instruct777https://huggingface.co/unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit, using unsloth library888https://github.com/unslothai/unsloth.
For the ‘after’ text generation subtask from the iterative self-writing experiment, we used the following hyperparameter setups for the fine-tuned Llama-8B-Instruct: (1) only one epoch of training; (2) a weight decay of 0.01; (3) warm-up steps of 10; (4) learning rate of ; and (5) AdamW 8-bit optimizer. Due to computational constraints, we were able to run only one epoch for fine-tuning Llama models on our ScholaWrite dataset. Also, it took approximately 12 hours on one NVIDIA L40s.
GPT-4o
We used the GPT-4o-2024-08-06 version. The iterative writing with 100 iterations took approximately 1 hour on each seed.
G.2 Details About Finetuning Process
The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. For fine-tuning, we used QLoRA (Dettmers et al., 2024) to optimize all linear modules of a 4-bit quantized model while maintaining a small memory footprint. Additionally, the embed_tokens and lm_head modules were set as trainable and saved in the final checkpoint. To focus training on the response portion (the “after” text), we used the train_on_response_only function, masking the task instructions, verbalizer, and “before” text with -100s. This ensures the model learns to generate the "after" text without being influenced by instructional input. The model was trained for one epoch with a batch size of 1, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
G.3 Prompt Templates
G.3.1 Generation Prompt for Llama-8B-Zero models
## Input: before_text ”
G.3.2 Generation Prompt for Llama-8B-Zero models
## Input: before_text ”
G.3.3 Generation Prompt for GPT-4o